# Hyperfocal subalgebras of hyperfocal abelian Frobenius blocks

### v1

@@ 2 Preliminaries
Denote $\mbox{$\mathcal{K}$}\mathop{\otimes}\limits_{\mathcal{O}}C$ and $k\mathop{\otimes}\limits_{\mathcal{O}}C$ by $\hat{C}$ and $\bar{C}$, respectively. We have two Grothendieck groups $R_{\mathcal{K}}(C)$ and $R_{k}(C)$ associated with these two algebras $\hat{C}$ and $\bar{C}$, respectively. This means that $R_{\mathcal{K}}(C)$ and $R_{k}(C)$ are the Grothendieck groups of finite dimensional $\hat{C}-$modules and $\bar{C}$-modules, respectively. Furthermore, denote by $\mathrm{Pr}_{\mathcal{O}}(C)$ the subgroup of $R_{\mathcal{K}}(C)$ generated by the images of modules of the form $\hat{C}\mathfrak{i}$ for some idempotent $\mathfrak{i}$ of $C$. For any finite dimensional module $U$ over $\hat{C}-$modules or $\bar{C}$-modules, we denote by $[U]$ the image of $U$ in $R_{\mathcal{K}}(C)$ or $R_{k}(C)$, respectively. It is clear that the set of images of simple $\hat{C}-$modules, denoted by $\mathrm{Irr}_{\mathcal{K}}(C)$, is a $\mathbb{Z}$-basis of $R_{\mathcal{K}}(C)$. Set $\mathrm{k}(C)=|\mathrm{Irr}_{\mathcal{K}}(C)|$. Similarly, the set of images of simple $\bar{C}$-modules, denoted by $\mathrm{Irr}_{k}(C)$, is a $\mathbb{Z}$-basis of $R_{k}(C)$. Set $l(C)=|\mathrm{Irr}_{k}(C)|$.
@@
Furthermore, for any finite dimensional $\bar{C}$-module $X$, there exists a finitely generated $C$-module $Y$ which is free of finite rank as an $\mathcal{O}$-module such that $\mathcal{K}\mathop{\otimes}\limits_{\mathcal{O}}Y$ is isomorphic to $X$. This induces the so-called decomposition map $d_{C}$ from $R_{\mathcal{K}}(C)$ to $R_{k}(C)$, sending $[X]$ to $k\mathop{\otimes}\limits_{\mathcal{O}}Y]$. Its kernel is denoted by $L^{0}(C)$. The nonnegative integers occurring in the matrix of $d_{C}$ with respect to the canonical bases $\mathrm{Irr}_{\mathcal{K}}(C)$ and $\mathrm{Irr}_{k}(C)$ are called decomposition numbers. It is well known that there is a bilinear form $\langle,\rangle_{C}$ on $R_{\mathcal{K}}(C)$ such that $\langle[U],[V]\rangle_{C}=\mathrm{dim}_{\mathcal{K}}(U,V)$ for any two finite dimensional $\hat{C}-$modules $U$ and $V$. Set $L^{0}(C)^{\bot}$ to be the subgroup of $R_{\mathcal{K}}(\mathbb{A})$ consisting of elements orthogonal to every elements of $L^{\circ}(C)$ with respect to this bilinear form.
@@
We fix a finite group $H$ in the remaining paragraphs of this section. An $\mathcal{O}$-algebra $\mathbb{B}$ is called an $H$-algebra if there is a group homomorphism $H\longrightarrow\mathrm{Aut}_{\mathcal{O}}(\mathbb{B})$. If this group homomorphism is induced by a group homomorphism $\theta:H\longrightarrow\mathbb{B}^{*}$, the $H$-algebra $\mathbb{B}$ is called an $H$-interior. Let $\mathbb{B}$ be an $H$-interior algebra. Clearly, $\mathbb{B}$ becomes an $\mbox{$\mathcal{O}$}H$-$\mbox{$\mathcal{O}$}H$-bimodule via this group homomorphism. We simply write $\theta(h^{-1})(a)\theta(h)$ as $a^{h}$ and $\varphi(x)a\varphi(y)$ as xayxay for any $h,x,y\in H$ any $a\in\mathbb{B}$. Let $\mathbb{B}^{\prime}$ be another $H$-interior algebra. An $\mathcal{O}$-algebra homomorphism $f:\mathbb{B}\rightarrow\mathbb{B}^{\prime}$ is called an $H$-interior algebra homomorphism if $f$ preserves the $H$-interior algebra structures on $\mathbb{B}$ and $\mathbb{B}^{\prime}$. Here, $f$ needs not be unitary. We say it is an embedding if $\text{Ker}(f)=\{0\}$ and $\text{Im}(f)=f(1)\mathbb{B}^{\prime}f(1)$. For any subgroup $Z$ of $H$, denote by $\mathbb{B}^{Z}$ the subalgebra of all elements of $\mathbb{B}$ fixed by the $Z$-conjugation. The $H$-interior algebra $\mathbb{B}$ is called primitive if $\mathbb{B}^{H}$ is a local $\mathcal{O}$-algebra.
@@
A point $\alpha$ of $Z$ on $\mathbb{B}$ is a $(\mathbb{B}^{Z})^{*}$-conjugacy class of primitive idempotents in $\mathbb{B}^{Z}$. The pair $(Z,\alpha)$, denoted by $Z_{\alpha}$, is called a pointed group on $\mathbb{B}$. We call a $Z$-interior algebra $\mathbb{B}_{\alpha}$ an embedded algebra of $Z_{\alpha}$ if there is an embedding of $Z$-interior algebras $f_{\alpha}:\mathbb{B}_{\alpha}\longrightarrow\mathbb{B}$ such that $f_{\alpha}(1_{\mathbb{B}_{\alpha}})\in\alpha$. By [20, 1.6], embedded algebras of $Z_{\alpha}$ are unique up to isomorphism. A typical choice of embedded algebras of $Z_{\alpha}$ is $a\mathbb{B}a$ and then the embedding $f_{\alpha}:a\mathbb{B}a\longrightarrow\mathbb{B}$ is just induced by inclusion. Here, $a$ is an element of $\alpha$ and $\mathbb{B}_{\alpha}$ is a $Z$-interior algebra with the group homomorphism $Z\rightarrow(\mathbb{B}_{\alpha})^{*},z\mapsto za$. We denote its image by $Za$. For two pointed groups $Z_{\alpha}$ and $Z^{\prime}_{\alpha^{\prime}}$ on $\mathbb{B}$, we say that $Z_{\alpha}$ is contained in $Z^{\prime}_{\alpha^{\prime}}$ if $Z\leq Z^{\prime}$ and there are $e\in\alpha$ and $e^{\prime}\in\alpha^{\prime}$ such that $ee^{\prime}=e=e^{\prime}e$. It is clear that the subgroup $N_{H}(Z)$ acts on the set $\{Z_{\alpha}~|~\alpha~\mathrm{runs~over~all~points~of~}Z~\mathrm{on~}\mathbb{B}\}$ by conjugation. Denote by $N_{H}(Z_{\alpha})$ the stabilizer of the pointed group $Z_{\alpha}$ under this action.
@@
For any subgroup $K$ of $Z$, we have the relative trace map $\mathrm{Tr}_{K}^{Z}:\mathbb{B}^{K}\rightarrow\mathbb{B}^{Z},a\mapsto\sum\limits_{z}a^{z}$, where $z$ runs over a set of representatives of the right cosets of $K$ in $Z$. Denote by $\mathbb{B}_{K}^{Z}$ the image of the map $\mathrm{Tr}_{K}^{Z}$. The sum $\sum\limits_{K<Z}\mathbb{B}_{K}^{Z}+J(\mathcal{O})\mathbb{B}^{Z}$ is an ideal of $\mathbb{B}^{Z}$. Denote by $\mathbb{B}(Z)$ the quotient of $\mathbb{B}^{Z}$ by this ideal and by $\mathrm{Br}_{Z}^{\mathbb{B}}$ the canonical homomorphism $\mathbb{B}^{Z}\rightarrow\mathbb{B}(Z)$. The quotient $\mathbb{B}(Z)$ is called the Brauer quotient of $\mathbb{B}$ at $Z$ and the associated homomorphism $\mathrm{Br}_{Z}^{\mathbb{B}}$ is called the Brauer homomorphism. If no confusion arises, we denote $\mathrm{Br}_{Z}^{\mathbb{B}}$ by $\mathrm{Br}_{Z}$. By [29, Lemma 11.7], $\mathbb{B}(Z)\neq 0$ only if $Z$ is a $p$-group. When $\mathbb{B}$ is the group algebra $\mathcal{O}L$ for some finite group $L$, the inclusion $\mathcal{O}C_{L}(Z)\subseteq(\mathcal{O}L)^{Z}$ induces a $k$-algebra isomorphism $(\mathcal{O}L)(Z)\cong kC_{L}(Z)$. We identify these two $k$-algebras through the isomorphism. More generally, there is a similar construction for any $\mathcal{O}Z$-module $M$. Set $M^{K}=\{m\in M\mid x\cdot m=m,\forall x\in M\}$. We can define a relative trace map $\mathrm{Tr}_{K}^{Z}:M^{K}\rightarrow M^{Z}$ by a similar way and denote its image by $M_{K}^{Z}$. Similarly, we set $M(Z)$ to be the quotient $\mathcal{O}$-module $M^{Z}/(\sum\limits_{K<Z}M_{K}^{Z}+J(\mathcal{O})M^{Z})$ and also denote the canonical homomorphism $M^{Z}\rightarrow M(Z)$ by $\mathrm{Br}_{Z}^{M}$, or $\mathrm{Br}_{Z}$ for simplicity if no confusion arises.
@@
Let $R_{\gamma}$ be a local point on $\mathbb{B}$. Take an element $i\in\gamma$ and set $\mathbb{B}_{\gamma}=i\mathbb{B}i$. Denote the subgroup $\{a\in\mathbb{B}_{\gamma}^{*}\mid{{}^{a}}(Ri)=Ri\}$ of $\mathbb{B}_{\gamma}^{*}$ by $N_{\mathbb{B}_{\gamma}}(R)$. It has three normal subgroups $Ri$ and $(\mathbb{B}_{\gamma}^{R})^{*}$ and $1+J(\mathbb{B}_{\gamma}^{R})$. Denote the two quotient groups $N_{\mathbb{B}_{\gamma}}(R)/Ri\cdot(1+J(\mathbb{B}_{\gamma}^{R}))$ and $N_{\mathbb{B}_{\gamma}}(R)/Ri\cdot(\mathbb{B}_{\gamma}^{R})^{*}$ by $\bar{F}_{\mathbb{B}}(R)$ and $F_{\mathbb{B}}(R)$, respectively. It is well-known that $(\mathbb{B}_{\gamma}^{R})^{*}$ is isomorphic to $k^{*}\times(1+J(\mathbb{B}_{\gamma}^{R}))$. We can get the following short exact sequence
${1\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${k^{*}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\bar{F}_{\mathbb{B}}(R)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${F_{\mathbb{B}}(R)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${1.}$
When $\mathbb{B}$ is the group algebra $\mbox{$\mathcal{O}$}L$ for some finite group $L$, by [20, Theorem 3.1], there is a group isomorphism between $N_{L}(R_{\gamma})/RC_{L}(R)$ and $F_{\mathcal{O}L}(R)$ which preserves the actions on $R$ up to inner automorphisms.
@@
For any element $x$ in $H$ and any subgroup $K$ of $H$, we denote by $[x,K]$ the subgroup of $H$ generated by elements $xux^{-1}u^{-1}$, where $u$ runs over $K$. Let $\mathfrak{b}$ be a block of $H$. Then $H_{\{\mathfrak{b}\}}$ is a pointed group on the $H$-interior algebra $\mathcal{O}H$. Let $Q_{\delta}$ be a defect pointed group of the block $\mathfrak{b}$ and $A$ a source algebra of $\mathcal{O}H\mathfrak{b}$. Then $Q$ can act on $A$ through the conjugation. Recall that the hyperfocal subgroup $\mathfrak{hyp}(Q_{\delta})$ of $Q_{\delta}$ is the subgroup of $Q$ generated by the subgroups $[x,T]$ where $T_{\epsilon}$ runs over the set of local pointed groups on $\mathcal{O}H$ such that $T_{\epsilon}\subseteq Q_{\delta}$, and $x$ over the set of $p^{\prime}$-elements of $N_{H}(T_{\epsilon})$ (see [25, 1.7.2]). Sometimes, we also call $\mathfrak{hyp}(Q_{\delta})$ a hyperfocal subgroup of $Q$ for simplicity. It is easy to see that $\mathfrak{hyp}(Q_{\delta})$ is a normal subgroup of $Q$. Denote by $\mathcal{S}$ a set of representatives for $Q/\mathfrak{hyp}(Q_{\delta})$ in $Q$. By [25, Theorem 1.8], there is a $Q$-stable $\mathcal{O}$-subalgebra $\mathbb{A}$ of $A$, unique up to $(A^{Q})^{*}$-conjugation, containing the image of $\mathfrak{hyp}(Q_{\delta})$ and fulfilling $A=\mathop{\bigoplus}\limits_{v\in\mathcal{S}}\mathbb{A}\cdot v$. It is clear that $\mathbb{A}$ is an $\mathfrak{hyp}(Q_{\delta})$-interior algebra inherited from the $Q$-interior algebra structure of $A$. The $\mathfrak{hyp}(Q_{\delta})$-interior algebra $\mathbb{A}$ is called a hyperfocal subalgebra of $A$. Sometimes, we also call it a hyperfocal subalgebra of the block $\mathfrak{b}$.
In the following paragraph, we will collect some notation and definition about endopermutation modules over a $p$-group which is introduced by Dade [4, 5].
@@
Let $P$ be a $p$-group. An $\mathcal{O}P$-module $V$ is called an endopermutation $\mathcal{O}P$-module if the endomorphism algebra $\mathrm{End}_{\mathcal{O}}(V)$ has a $P$-stable $\mathcal{O}$-basis under the $P$-conjugation action. Let $V$ be an indecomposable endopermutation $\mbox{$\mathcal{O}$}P$-module with vertex $P$, which means that $S(P)\neq 0$. Here, $S=\mathrm{End}_{\mathcal{O}}(V)$, which is called a Dade $P$-interior algebra. Then by [17, Proposition 7.3.7], for any subgroup $Q$ of $P$, there is up to isomorphism a unique endopermutation $k(N_{P}(Q)/Q)$-module $W_{Q}$ such that $S(Q)\cong\mathrm{End}_{k}(W_{Q})$ as $N_{P}(Q)/Q$-algebras. It is clear that $W_{Q}$ is an indecomposable endopermutation $k(N_{P}(Q)/Q)$-module. Moreover, let $R$ be a subgroup of $P$ such that $Q$ is normal in $R$. We denote by $\emph{Defres}_{R/Q}^{P}(V)$ the endopermutation $k(R/Q)$-module $\mathrm{Res}_{R/Q}^{N_{P}(Q)/Q}(W_{Q})$. In particular, when $Q$ is normal in $P$, $\mathrm{Defres}_{P/Q}^{P}(V)$ is just the indecomposable endopermutation $kP/Q$-module $W_{Q}$.
@@
Finally, let us recall the definitions of twisted group algebra and relative separability. Let $\alpha\in Z^{2}(H;k^{*})$ be a $2$-cocycle of $H$. Since $k$ is perfect, we have a canonical isomorphism $\mbox{$\mathcal{O}$}^{*}\cong k^{*}\times(1+J(\mbox{$\mathcal{O}$}))$ (see [28]). So we can view $k^{*}$ as a subgroup of $\mbox{$\mathcal{O}$}^{*}$. The twisted group algebra of $H$ by $\alpha$, denoted by $\mbox{$\mathcal{O}$}_{\alpha}H$, is the $\mathcal{O}$-algebra with an $\mathcal{O}$-basis consisting of all elements of $H$ and the multiplication induced by the equality $\lambda_{x}x\cdot\lambda_{y}y=(\lambda_{x}\lambda_{y}\alpha(x,y))xy$ for all $x,y\in H$ and all $\lambda_{x},\lambda_{y}\in\mbox{$\mathcal{O}$}$. An $H$-interior algebra $\mathbb{B}$ is called relatively $\mbox{$\mathcal{O}$}H$-separable if $\mathbb{B}$ is isomorphic to a direct summand of $\mathbb{B}\mathop{\otimes}\limits_{\mathcal{O}H}\mathbb{B}$ as $\mathbb{B}$-$\mathbb{B}$-bimodules (see [16, §4.8]). This is equivalent to requiring that the canonical $\mathbb{B}$-$\mathbb{B}$-bimodule homomorphism $\mathbb{B}\mathop{\otimes}\limits_{\mathcal{O}H}\mathbb{B}\longrightarrow\mathbb{B}$ induced by multiplication in $\mathbb{B}$ splits by [16, Theorem 2.6.10].

### v2

@@ 2 Preliminaries
Denote $\mbox{$\mathcal{K}$}\mathop{\otimes}\limits_{\mathcal{O}}C$ and $k\mathop{\otimes}\limits_{\mathcal{O}}C$ by $\hat{C}$ and $\bar{C}$, respectively. We have two Grothendieck groups $R_{\mathcal{K}}(C)$ and $R_{k}(C)$ associated with these two algebras $\hat{C}$ and $\bar{C}$, respectively. This means that $R_{\mathcal{K}}(C)$ and $R_{k}(C)$ are the Grothendieck groups of finite dimensional $\hat{C}-$modules and $\bar{C}$-modules, respectively. Furthermore, denote by $\mathrm{Pr}_{\mathcal{O}}(C)$ the subgroup of $R_{\mathcal{K}}(C)$ generated by the images of modules of the form $\hat{C}\mathfrak{i}$ for some idempotent $\mathfrak{i}$ of $C$. For any finite dimensional module $U$ over $\hat{C}$ or $\bar{C}$, we denote by $[U]$ the image of $U$ in $R_{\mathcal{K}}(C)$ or $R_{k}(C)$, respectively. It is clear that the set of images of simple $\hat{C}-$modules, denoted by $\mathrm{Irr}_{\mathcal{K}}(C)$, is a $\mathbb{Z}$-basis of $R_{\mathcal{K}}(C)$. When $C$ is a group algebra $\mbox{$\mathcal{O}$}H$ for some finite group $H$, we identify ${\rm Irr}_{\mathcal{K}}(C)$ with ${\rm Irr}(H)$ if no confusion arises throughout this paper. Set $\mathrm{k}(C)=|\mathrm{Irr}_{\mathcal{K}}(C)|$. Similarly, the set of images of simple $\bar{C}$-modules, denoted by $\mathrm{Irr}_{k}(C)$, is a $\mathbb{Z}$-basis of $R_{k}(C)$. Set $l(C)=|\mathrm{Irr}_{k}(C)|$.
@@
Furthermore, for any finite dimensional $\hat{C}-$module $X$, there exists a finitely generated $C$-module $Y$ which is free of finite rank as an $\mathcal{O}$-module such that $\mathcal{K}\mathop{\otimes}\limits_{\mathcal{O}}Y$ is isomorphic to $X$ as $\hat{C}-$modules. This induces the so-called decomposition map $d_{C}$ from $R_{\mathcal{K}}(C)$ to $R_{k}(C)$, sending $[X]$ to $k\mathop{\otimes}\limits_{\mathcal{O}}Y]$. Its kernel is denoted by $L^{0}(C)$. The nonnegative integers occurring in the matrix of $d_{C}$ with respect to the canonical bases $\mathrm{Irr}_{\mathcal{K}}(C)$ and $\mathrm{Irr}_{k}(C)$ are called decomposition numbers. It is well-known that there is a bilinear form $\langle~,~\rangle_{C}$ on $R_{\mathcal{K}}(C)$ such that $\langle[U],[V]\rangle_{C}=\mathrm{dim}_{\mathcal{K}}{\rm Hom}_{\hat{C}}(U,V)$ for any two finite dimensional $\hat{C}-$modules $U$ and $V$. Set $L^{0}(C)^{\bot}$ to be the subgroup of $R_{\mathcal{K}}(C)$ consisting of elements orthogonal to every elements of $L^{0}(C)$ with respect to this bilinear form.
@@
We fix a finite group $H$ in the remaining paragraphs of this section. An $\mathcal{O}$-algebra $\mathbb{B}$ is called an $H$-algebra if there is a group homomorphism $H\longrightarrow\mathrm{Aut}_{\mathcal{O}}(\mathbb{B})$. If this group homomorphism is induced by a group homomorphism $\theta:H\longrightarrow\mathbb{B}^{\times}$, the $H$-algebra $\mathbb{B}$ is called an $H$-interior algebra. Let $\mathbb{B}$ be an $H$-interior algebra. Clearly, $\mathbb{B}$ becomes an $\mbox{$\mathcal{O}$}H$-$\mbox{$\mathcal{O}$}H$-bimodule via this group homomorphism. We simply write $\theta(h^{-1})(a)\theta(h)$ as $a^{h}$ and $\theta(x)a\theta(y)$ as xayxay for any $h,x,y\in H$ and any $a\in\mathbb{B}$. Let $\mathbb{B}^{\prime}$ be another $H$-interior algebra. An $\mathcal{O}$-algebra homomorphism $f:\mathbb{B}\longrightarrow\mathbb{B}^{\prime}$ is called an $H$-interior algebra homomorphism if $f$ preserves the $H$-interior algebra structures on $\mathbb{B}$ and $\mathbb{B}^{\prime}$. Here, $f$ needs not be unitary. We say it is an embedding if $\text{Ker}(f)=\{0\}$ and $\text{Im}(f)=f(1)\mathbb{B}^{\prime}f(1)$. For any subgroup $Z$ of $H$, denote by $\mathbb{B}^{Z}$ the subalgebra of all elements of $\mathbb{B}$ fixed by the $Z$-conjugation. The $H$-interior algebra $\mathbb{B}$ is called primitive if $\mathbb{B}^{H}$ is a local $\mathcal{O}$-algebra.
@@
A point $\alpha$ of $Z$ on $\mathbb{B}$ is a $(\mathbb{B}^{Z})^{\times}$-conjugacy class of primitive idempotents in $\mathbb{B}^{Z}$. The pair $(Z,\alpha)$, denoted by $Z_{\alpha}$, is called a pointed group of $Z$ on $\mathbb{B}$. We call a $Z$-interior algebra $\mathbb{B}_{\alpha}$ an embedded algebra of $Z_{\alpha}$ if there is an embedding of $Z$-interior algebras $f_{\alpha}:\mathbb{B}_{\alpha}\longrightarrow\mathbb{B}$ such that $f_{\alpha}(1_{\mathbb{B}_{\alpha}})\in\alpha$. By [20, 1.6], embedded algebras of $Z_{\alpha}$ are unique up to isomorphism. A typical choice of embedded algebras of $Z_{\alpha}$ is $a\mathbb{B}a$ and then the embedding $f_{\alpha}:a\mathbb{B}a\longrightarrow\mathbb{B}$ is just induced by inclusion. Here, $a$ is an element of $\alpha$ and $\mathbb{B}_{\alpha}$ is a $Z$-interior algebra with the group homomorphism $Z\rightarrow(\mathbb{B}_{\alpha})^{\times},z\mapsto za$. We denote its image by $Za$. For two pointed groups $Z_{\alpha}$ and $Z^{\prime}_{\alpha^{\prime}}$ on $\mathbb{B}$, we say that $Z_{\alpha}$ is contained in $Z^{\prime}_{\alpha^{\prime}}$ if $Z\leq Z^{\prime}$ and there are $e\in\alpha$ and $e^{\prime}\in\alpha^{\prime}$ such that $ee^{\prime}=e=e^{\prime}e$. It is clear that the subgroup $N_{H}(Z)$ acts on the set $\{Z_{\alpha}~|~\alpha~\mathrm{runs~over~all~points~of~}Z~\mathrm{on~}\mathbb{B}\}$ by conjugation. Denote by $N_{H}(Z_{\alpha})$ the stabilizer of the pointed group $Z_{\alpha}$ under this action.
@@
For any subgroup $K$ of $Z$, we have the relative trace map $\mathrm{Tr}_{K}^{Z}:\mathbb{B}^{K}\rightarrow\mathbb{B}^{Z},a\mapsto\sum\limits_{z}a^{z}$, where $z$ runs over a set of representatives of the right cosets of $K$ in $Z$. Denote by $\mathbb{B}_{K}^{Z}$ the image of the map $\mathrm{Tr}_{K}^{Z}$. The sum $\sum\limits_{K<Z}\mathbb{B}_{K}^{Z}+J(\mathcal{O})\mathbb{B}^{Z}$ is an ideal of $\mathbb{B}^{Z}$. Denote by $\mathbb{B}(Z)$ the quotient of $\mathbb{B}^{Z}$ by this ideal and by $\mathrm{Br}_{Z}^{\mathbb{B}}$ the canonical homomorphism $\mathbb{B}^{Z}\longrightarrow\mathbb{B}(Z)$. The quotient $\mathbb{B}(Z)$ is called the Brauer quotient of $\mathbb{B}$ at $Z$ and the associated homomorphism $\mathrm{Br}_{Z}^{\mathbb{B}}$ is called the Brauer homomorphism. If no confusion arises, we denote $\mathrm{Br}_{Z}^{\mathbb{B}}$ by $\mathrm{Br}_{Z}$. By [29, Lemma (11.7)], $\mathbb{B}(Z)\neq 0$ only if $Z$ is a $p$-group. When $\mathbb{B}$ is the group algebra $\mathcal{O}L$ for some finite group $L$, the inclusion $\mathcal{O}C_{L}(Z)\subseteq(\mathcal{O}L)^{Z}$ induces a $k$-algebra isomorphism $(\mathcal{O}L)(Z)\cong kC_{L}(Z)$. We identify these two $k$-algebras through the isomorphism. More generally, there is a similar construction for any $\mathcal{O}Z$-module $M$. Set $M^{K}=\{m\in M\mid x\cdot m=m,\forall x\in M\}$. We can define a relative trace map $\mathrm{Tr}_{K}^{Z}:M^{K}\longrightarrow M^{Z}$ by a similar way and denote its image by $M_{K}^{Z}$. Similarly, we set $M(Z)$ to be the quotient $\mathcal{O}$-module $M^{Z}/(\sum\limits_{K<Z}M_{K}^{Z}+J(\mathcal{O})M^{Z})$ and also denote the canonical homomorphism $M^{Z}\rightarrow M(Z)$ by $\mathrm{Br}_{Z}^{M}$, or $\mathrm{Br}_{Z}$ for simplicity if no confusion arises.
@@
Let $R_{\gamma}$ be a local point on $\mathbb{B}$. Take an element $i\in\gamma$ and set $\mathbb{B}_{\gamma}=i\mathbb{B}i$. Denote the subgroup $\{a\in\mathbb{B}_{\gamma}^{\times}\mid{{}^{a}}(Ri)=Ri\}$ of $\mathbb{B}_{\gamma}^{\times}$ by $N_{\mathbb{B}_{\gamma}}(R)$. It has three normal subgroups $Ri$ and $(\mathbb{B}_{\gamma}^{R})^{\times}$ and $1+J(\mathbb{B}_{\gamma}^{R})$. Denote the two quotient groups $N_{\mathbb{B}_{\gamma}}(R)/Ri\cdot(1+J(\mathbb{B}_{\gamma}^{R}))$ and $N_{\mathbb{B}_{\gamma}}(R)/Ri\cdot(\mathbb{B}_{\gamma}^{R})^{\times}$ by $\bar{F}_{\mathbb{B}}(R)$ and $F_{\mathbb{B}}(R)$, respectively. It is well-known that $(\mathbb{B}_{\gamma}^{R})^{\times}$ is isomorphic to $k^{\times}\times(1+J(\mathbb{B}_{\gamma}^{R}))$. We can get the following short exact sequence
${1\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${k^{\times}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\bar{F}_{\mathbb{B}}(R)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${F_{\mathbb{B}}(R)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${1.}$
When $\mathbb{B}$ is the group algebra $\mbox{$\mathcal{O}$}L$ for some finite group $L$, by [20, Theorem 3.1], there is a group isomorphism $\mathfrak{f}_{L,R}$ between $N_{L}(R_{\gamma})/RC_{L}(R)$ and $F_{\mathcal{O}L}(R)$, which sends $xRC_{L}(R)\in N_{L}(R_{\gamma})/RC_{L}(R)$ to $a_{x}(Ri\cdot(\mathbb{B}_{\gamma}^{R})^{\times})\in F_{\mathbb{B}}(R)$ such that $a_{x}$ can be chosen to satisfy $^{a_{x}}(ri)=({{}^{x}}r)i$ for any $r\in R$.
@@
For any element $x$ in $H$ and any subgroup $K$ of $H$, we denote by $[x,K]$ the subgroup of $H$ generated by elements $xux^{-1}u^{-1}$, where $u$ runs over $K$. Let $\mathfrak{b}$ be a block of $H$. Then $H_{\{\mathfrak{b}\}}$ is a pointed group on the $H$-interior algebra $\mathcal{O}H$. A defect pointed group of the pointed group $H_{\{\mathfrak{b}\}}$ is also called a defect pointed group of the block $\mathfrak{b}$. Let $Q_{\delta}$ be a defect pointed group of the block $\mathfrak{b}$ and $A$ a source algebra of $\mathcal{O}H\mathfrak{b}$. Then $Q$ can act on $A$ through the conjugation. Recall that the hyperfocal subgroup $\mathfrak{hyp}(Q_{\delta})$ of $Q_{\delta}$ is the subgroup of $Q$ generated by the subgroups $[x,R]$ where $R_{\epsilon}$ runs over the set of local pointed groups on $\mathcal{O}H$ such that $R_{\epsilon}\subseteq Q_{\delta}$, and $x$ over the set of $p^{\prime}$-elements of $N_{H}(R_{\epsilon})$ (see [25, 1.7.2]). Sometimes, we also call $\mathfrak{hyp}(Q_{\delta})$ a hyperfocal subgroup of $Q$ for simplicity. It is easy to see that $\mathfrak{hyp}(Q_{\delta})$ is a normal subgroup of $Q$. Denote by $\mathcal{S}$ a set of representatives for $Q/\mathfrak{hyp}(Q_{\delta})$ in $Q$. By [25, Theorem 1.8], there is a $Q$-stable $\mathcal{O}$-subalgebra $\mathbb{A}$ of $A$, unique up to $(A^{Q})^{\times}$-conjugation, containing the image of $\mathfrak{hyp}(Q_{\delta})$ and fulfilling $A=\mathop{\bigoplus}\limits_{v\in\mathcal{S}}\mathbb{A}v=\mathop{\bigoplus}\limits_{v\in\mathcal{S}}v\mathbb{A}$. It is clear that $\mathbb{A}$ is an $\mathfrak{hyp}(Q_{\delta})$-interior algebra inherited from the $Q$-interior algebra structure of $A$. The $\mathfrak{hyp}(Q_{\delta})$-interior algebra $\mathbb{A}$ is called a hyperfocal subalgebra of $A$. Sometimes, we also call it a hyperfocal subalgebra of the block $\mathfrak{b}$.
In the following paragraph, we will collect some notation and definitions about endopermutation modules over a $p$-group which is introduced by Dade [4, 5].
@@
Let $P$ be a $p$-group. An $\mathcal{O}P$-module $V$ is called an endopermutation $\mathcal{O}P$-module if the endomorphism algebra $\mathrm{End}_{\mathcal{O}}(V)$ has a $P$-stable $\mathcal{O}$-basis under the $P$-conjugation action. Let $V$ be an indecomposable endopermutation $\mbox{$\mathcal{O}$}P$-module with vertex $P$, which means that $S(P)\neq 0$. Here, $S=\mathrm{End}_{\mathcal{O}}(V)$, which is called a Dade $P$-interior algebra. Then by [16, Proposition 7.3.7], for any subgroup $Q$ of $P$, there is up to isomorphism a unique endopermutation $k(N_{P}(Q)/Q)$-module $W_{Q}$ such that $S(Q)\cong\mathrm{End}_{k}(W_{Q})$ as $N_{P}(Q)/Q$-algebras. It is clear that $W_{Q}$ is an indecomposable endopermutation $k(N_{P}(Q)/Q)$-module. Moreover, let $R$ be a subgroup of $P$ such that $Q$ is normal in $R$. We denote by $\emph{Defres}_{R/Q}^{P}(V)$ the endopermutation $k(R/Q)$-module $\mathrm{Res}_{R/Q}^{N_{P}(Q)/Q}(W_{Q})$. In particular, when $Q$ is normal in $P$, $\mathrm{Defres}_{P/Q}^{P}(V)$ is just the indecomposable endopermutation $k(P/Q)$-module $W_{Q}$.
@@
Let us recall the definitions of twisted group algebra and relative separability. Consider the trivial action of $H$ on $k^{\times}$. We denote by $Z^{2}(H;k^{\times})$ and $H^{2}(H;k^{\times})$ the set of all $2$-cocycles of $H$ with coefficients in $k^{\times}$ and the second cohomology group of $H$ with coefficients in $k^{\times}$, respectively. Let $\alpha\in Z^{2}(H;k^{\times})$ be a $2$-cocycle of $H$. Since $k$ is perfect, we have a canonical isomorphism $\mbox{$\mathcal{O}$}^{\times}\cong k^{\times}\times(1+J(\mbox{$\mathcal{O}$}))$ (see [28]). So we can view $k^{\times}$ as a subgroup of $\mbox{$\mathcal{O}$}^{\times}$. The twisted group algebra of $H$ by $\alpha$, denoted by $\mbox{$\mathcal{O}$}_{\alpha}H$, is the $\mathcal{O}$-algebra with an $\mathcal{O}$-basis consisting of all elements of $H$ and the multiplication induced by the equality $\lambda_{x}x\cdot\lambda_{y}y=(\lambda_{x}\lambda_{y}\alpha(x,y))xy$ for all $x,y\in H$ and all $\lambda_{x},\lambda_{y}\in\mbox{$\mathcal{O}$}$. Let $\mathbb{B}$ and $\mathbb{B}^{\prime}$ be two $\mathcal{O}$-algebras. Let $\varphi$ be an $\mathcal{O}$-algebra homomorphism from $\mathbb{B}^{\prime}$ to $\mathbb{B}$. Through this homomorphism $\varphi$, we can view $\mathbb{B}$ as a $\mathbb{B}^{\prime}$-$\mathbb{B}^{\prime}$-bimodule. The $\mathcal{O}$-algebra $\mathbb{B}$ is called relatively $\mathbb{B}^{\prime}$-separable if $\mathbb{B}$ is isomorphic to a direct summand of $\mathbb{B}\mathop{\otimes}\limits_{\mathbb{B}^{\prime}}\mathbb{B}$ as $\mathbb{B}$-$\mathbb{B}$-bimodules (see [15, §4.8]). This is equivalent to requiring that the canonical $\mathbb{B}$-$\mathbb{B}$-bimodule homomorphism $\mathbb{B}\mathop{\otimes}\limits_{\mathbb{B}^{\prime}}\mathbb{B}\longrightarrow\mathbb{B}$ induced by multiplication in $\mathbb{B}$ splits by [15, Theorem 2.6.10]. In the case where $\mathbb{B}^{\prime}$ is a subalgebra of $\mathbb{B}$ the homomorphism $\varphi$ is always taken to be the inclusion map unless otherwise specified.